# Token Spammers, Rug Pulls, and SniperBots: An Analysis of the Ecosystem of Tokens in Ethereum and the Binance Smart Chain (BNB)

### v2

@@ 6.1 The Token Spammers
In this subsection we put under the lens the token spammers. Here are a few interesting observations.
@@
Token spammers are more prevalent in BSC. Although the distribution of the number of tokens created by addresses in Ethereum and BSC is almost identical (Figure 3), the absolute numbers are different. Indeed, in terms of raw numbers, we find that BSC has almost four times more token spammers than Ethereum (4,231 vs 1,329), and the spammers of BSC create almost four times more tokens in BSC than in Ethereum (262,023 vs. 67,838).
@@
Token spammers create tokens mainly with contract creation transaction. As mentioned in Section 2, tokens can be created in two ways: By sending a contract creation transaction or by sending a transaction to a smart contract that generates the token. We find that 94.8% of the tokens in BSC and 82.3% of the tokens in Ethereum are created directly by sending a contract creation transaction. Focusing on the small fraction created by sending a transaction to a smart contract, we find that token spammers create 13,794 tokens using 3,530 different contracts in BSC and 12,010 tokens using 1,237 different contracts in Ethereum.
@@
Token spammers create short lifetime tokens. As we have seen, a significant fraction of tokens have a lifetime shorter than 1 day. Investigating the relationship between token spammers and 1-day tokens, we discover that most of the tokens created by the spammers have a lifetime shorter than one day. The spammers created 170,768 1-day tokens out of 262,023 (65.1%) and 40,552 1-day tokens out of 67,838 (59.8%) respectively, in BSC and Ethereum.
7 Malicious Activity
@@
The top token spammer creates 17,936 tokens in the timeframe of our analysis. If we look at the name of these tokens, we find that almost all of them have the same name (the tokens have only 76 unique names), with the most used being ’Pornhub’ with 605 occurrences. The median lifetime of these tokens is extremely small: 45 mins. Lastly, almost all of the tokens (99.7%) created by this address have a liquidity pool. We now detail in depth the life of one of the tokens.
@@
The anatomy of an operation. We focus on OnlyFans3330xe8b6f08841d668605343A63144D76ff2dE9A1199, a token created by the top token spammer on block 8090747 (2021-06-07 01:40:34 PM UTC) by issuing a contract creation transaction. This token has a supply of $10^{32}$ and its symbol is the Unicode U+1F48B, the emoji of a kiss mark, followed by the string "OnlyFans". On block 8090751 (2021-06-07 01:40:46 PM UTC), after 4 blocks from its creation, the token spammer creates a liquidity pool that contains the pair (OnlyFans, WrappedBNB) and adds a liquidity of 20 Wrapped BNB (almost \$7,180 at the moment of the operation) and 44 trillions of OnlyFans tokens.
@@
After just 6 seconds, on block 8090753 (2021-06-07 01:40:52 PM UTC), an address swaps 4 billion OnlyFans for 0.002 Wrapped BNB (\$0.718). That operation is followed by other 12 addresses—12 different swaps—for a total buy of $5.2*10^{12}$ OnlyFans for 2.67 Wrapped BNB (\$958). After 2 hours from the creation of the token, at block 8093101 (2021-06-07 03:38:55 PM UTC), the token spammer removes all the liquidity from the liquidity pools leaving it drained. Since the 12 addresses added Wrapped BNB into the pool by buying OnlyFans, the token spammer collects 22.67 Wrapped BNB and has a profit of 2.67 Wrapped BNB (\$958).
These kind of operations are quite frequent and are commonly called rug pulls or exit scams [43, 37].
7.1 The Rug Pull
We develop an approach to systematically find operations following the same pattern we explored in the case study. First of all, we formalize the kind of activity we are interested in:
@@ 7.1 The Rug Pull
7.1.1 An Improved Version of the Fraud
The rug pull described above is the simple version of the fraud, in which the scammer never interact with the liquidity pools until he removes the liquidity. To attract more investors, the scammer can manipulate some statistics of the liquidity pool, such as the number of swaps, the trading volume, or the price. The scammer can usewash-trading [13], a well-known market manipulation. In this case, the creator of the pool tries to create the impression that the liquidity pool is active, faking the trading volume by repeatedly buying and selling tokens. The goal is to increase the liquidity pool’s trading volume rather than the scam token’s price.
Another way that scammers have to drum up the attention of investors is to inflate the price by buying the scam token gradually. This action is without risks for the scammer until the liquidity pool has no new investors. Indeed, having the total of the LP-tokens, he can retrieve all the new capital injected when he removes all the liquidity. In the following, we refer to these operations as pump operations.
The scammer can also hedge his gains—eliminating the risk of an unrealized profit while the liquidity pool is still active. The scammer usually does not put the scam token in the pool but maintains a reserve in its wallet. When investors start to buy the scam token, raising the asset’s price, the scammer can gradually sell the owned token, taking profit from the operation. In the following, we refer to these operations as hedge operations. Of course, the scammer can use one or a combination of these practices.
7.2 Looking for Rug Pulls
We leverage our datasets to identify exit scam rug pulls systematically. Since we saw a considerable number of 1-day tokens, and most of them are created serially, we narrow our investigation to the 414,936 in BSC and 26,817 in Ethereum 1-day tokens with a liquidity pool. We analyze all the Events emitted by the liquidity pools, looking for all the liquidity pools that emitted only 1 Mint event and only 1 Burn event in which the address that performs the transaction burns at least 99% of the minted LP-tokens (we don’t use 100% since a small fraction of tokens might be stuck in the wallet due to rounding).
7.2.1 Estimating the Gains of the Fraud
The simple operation, where the scammer does not swap in his liquidity pool, can be carried out by performing just four transactions: A transaction that creates the token, one that creates the liquidity pool, one that adds the liquidity, and finally, the last transaction to remove the liquidity. These transactions can be performed individually, or they can be aggregated by leveraging a smart contract. Of course, we consider both of the cases when computing the fees. If the scammer performs swaps on the liquidity pool, we also consider the transaction fees paid for each swap.
@@ 7.2.1 Estimating the Gains of the Fraud
The formula can be split into two components. The first part computes the gain in the case of the simple operation, the case that we explored in the Case Study and formalized in Section 7.1. The second formula takes into account the more refined version of the fraud that we described in 7.1.1, where the creator of the liquidity pool manipulates it by performing swaps operations. In this case, we remove from the gain $M_{in}$, that is, the amount of tokens that the manipulator artificially adds to the liquidity. We also add to the gain $T_{out}$, the quantity of tokens that the manipulator removes from the liquidity pool before the final removal of the liquidity. Finally, we remove from the gains the fees used to perform the swaps operations ($fee_{swap}$).
@@ 7.3 Results
After processing our data, we discover that an incredibly high number of liquidity pools are actually rug pulls. In BSC, 266,340 out of 414,936 (65.6%) of the considered liquidity pools have an exit scam pattern, while 21,594 out of 26,817 (81.1%) in Ethereum. This result shows that scammers use most of the 1-day tokens as disposable to carry out rug pulls.
These operations are arranged by 116,516 different addresses in BSC and 16,539 different addresses in Ethereum. As we can expect from the previous analyses, most of the token spammers that operate in BSC are linked to this kind of activity. Indeed, in BSC, 3,179 out of 4,231 (75%) token spammers performed at least one rug pull. Instead, in Ethereum, there are only 101 token spammers (7.6%) that have been involved in this activity. We find 114 addresses that perform more than 100 exit scams in BSC, accounting for 19.1% of the operations, with the most active performing 16,102 operations. Instead, in Ethereum, we find only two addresses performing more than 100 operations. Interestingly, combining the information in the BSC and Ethereum datasets, we find that two token spammers operated on both blockchains with the same address. They perform 78 exit scams in Ethereum and 194 in BSC.
Looking at the liquidity pools, we find that BNB (97.8% of the cases) is the token paired the most with the scam token. It is followed by USDT (0.67%) and BUSD (0.15%), two stable coins pegged to the USD. Instead, ETH is paired with all the scam tokens in all the liquidity pools with an exit scam in Ethereum. As the next step, we want to estimate the number of potential users that fall prey to such activities. To do so, we exclude the addresses that swap into liquidity pools they have created themselves from this analysis. We collect 256,967 different addresses in BSC and 59,043 in Ethereum that interact with at least one liquidity pool with an exit scam pattern. These addresses performed 2,903,021 swaps on the considered liquidity pools in BSC and 317,305 in Ethereum.
We divide the swaps into buy (scam token) and sell operations. As we can expect, given the anatomy of the fraud, we find that most of the operations are buy operations. More in details, in BSC 2,286,056 (78.7%) are buy operations and 616,966 (21.3%) sell operations. In Ethereum, we find a similar pattern, with 254,061 (80.1%) buy operations and 63,196 (19.9%) sell operations.
As final metric, we compute the average value of the swaps performed by the users. The average amount of swaps is almost identical for buy and sell operations on both the blockchains, with 0.01 BNB in BSC and 0.19 ETH in Ethereum. Interestingly, we notice a considerable difference in the average swap amount between the two blockchains. Indeed, the average swap is approximately \$3 in BSC and \$360 in Ethereum, considering the current price of BNB and ETH44407/06/2022.
@@ 7.3.1 The Gains
Before computing the gains of the scammers, we calculate the average price a scammer has to invest in arranging the fraud. If the scammer does not perform any swap into the liquidity pool, the cost of the operation is on average 0.03 BNB in the case of BSC and 0.2 ETH for the Ethereum blockchain. Thus, the investment needed to perform such operations is low, even if it could vary substantially when the blockchains are overloaded. For instance, we found some operations among our set of rug pulls that reached the cost of 1.1 BNB or even 3.3 ETH. The base cost to arrange the fraud is interesting because it represents a bound to the loss that the scammers have to afford for each operation.
@@
We leverage our datasets to compute the gain of the operation using Formula 1. We describe the 266,340 operations on BSC and the 21,594 in Ethereum in terms of successful and unsuccessful operations based on the operation’s net gain. In particular, we define an operation as successful if the net gain is strictly positive.
@@
Successful operations. Among the liquidity pools with an exit scam pattern, there are 104,404 (39.1%) operations in BSC and 13,368 (61.9%) in Ethereum closed with a profit for the scammer. A possible reason for the higher success rate of the rug pull in Ethereum could be that, as we saw, on average, users tend to invest more money. Indeed, on average, attracting only one investor is enough to cover the operation’s cost. To investigate what can affect the gains, we combine information on gains with those of the manipulation. When the creator of the liquidity pool does not perform any manipulation, the net gain is on average 0.11 BNB in BSC and 1.34 ETH in Ethereum. Operations carried out on liquidity pools that suffer wash-trading activity have an average gain of 0.25 BNB in BSC and 12 ETH in Ethereum, which is considerably higher than the previous case. Instead, we notice a negligible increase in gains in the case of pump operations with respect to the gains obtained by the liquidity pools without manipulations. Moreover, we find that both kinds of manipulation have no impact on the success rate. This show that operations that have wash trading are generally more profitable. However, there is a shortcoming. Indeed the scammer have to perform several swaps, increasing its cost and the loss in case of an unsuccessful operation.
@@
Unsuccessful operations. There are 161,936 (60.9%) liquidity pools in BSC and 8,226 (38.1%) in Ethereum, for which the scammer does not cover the transaction fees with the operations. For the 13% (21,122) of these liquidity pools of BSC and the 18.3% (1,506) of Ethereum, we notice that the operations were unsuccessful because nobody swapped into the liquidity pools. Considering the results we obtained, we can conjecture that the aim of the scammers is not to be successful every time but to arrange frauds serially and take profit in the long run. Indeed, the loss of an unsuccessful operation is minimal, and a streak of operations closed in loss can be covered with a single profitable operation.
7.3.2 The Tokens’ Names
Table 4: Tokens names most frequently used in exit scams.
BNB Smart Chain Ethereum
Name # of tokens Name # of tokens
Pornhub 1023 Hyve.works 50
Galaxy 588 Deriswap 32
Seedswap 502 Shibaswap 28
Lionswap 429 Apple core finance 17
Eco.finance 421 X20.finance 16
Spacex 419 Yield farm rice 15
Onlyfans 398 The sandbox 14

### v3

@@ 6 Token spammers
Token spammers are more prevalent in BSC. Although the distribution of the number of tokens created by addresses in Ethereum and BSC is almost identical (Fig. 3), the absolute numbers are different. Indeed, in terms of raw numbers, we find that BSC has almost four times more token spammers than Ethereum (4,231 vs. 1,329), and the spammers of BSC create almost four times more tokens in BSC than in Ethereum (262,023 vs. 67,838).
@@
Token spammers create tokens mainly with contract creation transactions. As mentioned in Section 2, tokens can be created in two ways: By sending a contract creation transaction or by sending a transaction to a smart contract that generates the token. We find that 94.8% of the tokens on BSC and 82.3% of the tokens on Ethereum are created directly by sending a contract creation transaction.
@@
Token spammers create short lifetime tokens. As we have seen, a significant amount of tokens have a lifetime shorter than one day. Investigating the relationship between token spammers and 1-day tokens, we discover that most of the tokens created by the spammers have a lifetime shorter than one day. The spammers created 170,768 1-day tokens out of 262,023 (65.1%) and 40,552 1-day tokens out of 67,869 (59.8%), respectively, in BSC and Ethereum.
7 The Anatomy of a Rug Pull
@@
The top token spammer creates 17,936 tokens in the timeframe of our analysis. If we look at the name of these tokens, we find that almost all of them have the same name (the tokens have only 76 unique names), with the most used being ’Pornhub’ with 605 occurrences. The median lifetime of these tokens is extremely small: 45 mins. Lastly, almost all of the tokens (99.7%) created by this address have a liquidity pool. We study the liquidity pools of these tokens and find out that they are used to perform an operation commonly known as rug pull [48, 43]. In the following, we report a detailed example of a rug pull operation carried out by this address.
@@
We focus on OnlyFans4440xe8b6f08841d668605343A63144D76ff2dE9A1199, a token created by the top token spammer on block 8090747 (2021-06-07 01:40:34 PM UTC) by issuing a contract creation transaction. On block 8090751 (2021-06-07 01:40:46 PM UTC), after 4 blocks from its creation, the token spammer creates a liquidity pool that contains the pair (OnlyFans, WrappedBNB) and adds a liquidity of 20 Wrapped BNB (almost \$7,180 at the moment of the operation) and 44 trillion of OnlyFans tokens.
@@
After just 6 seconds, on block 8090753, an address swaps 4 million OnlyFans for 0.002 Wrapped BNB (\$0.718). This operation is followed by 11 other swaps—performed by 11 different addresses—for a total buy of $5.1740396*10^{12}$ OnlyFans for 2.67 Wrapped BNB (\$958). After 2 hours from the creation of the token, at block 8093101 (2021-06-07 03:38:55 PM UTC), the token spammer removes all the liquidity from the liquidity pool, leaving it drained. Since the 12 addresses added Wrapped BNB into the pool by buying OnlyFans, the token spammer collects 22.67 Wrapped BNB and has a profit of 2.67 Wrapped BNB (\$958).
We can formalize these operations in the following way:
@@
An improved version of the operation. The rug pull described above is the simple version of the operation. However, to attract more investors, the attacker can manipulate some statistics of the liquidity pool. A well-known market manipulation that the attacker can use is wash-trading [12]. In this case, the creator of the pool tries to create the impression that the liquidity pool is active, faking the trading volume by repeatedly buying and selling tokens. Similarly, another way that attackers have to drum up the attention of investors is to inflate the price by buying the 1-day token gradually.
Finally, the attacker can also hedge his gains—eliminating the risk of an unrealized profit while the liquidity pool is still active. The attacker can maintain a reserve of tokens and, when investors start to buy the 1-day token, gradually sell the owned token, starting to take profit from the operation.
Clearly, rug pull operations can harm investors. However, we cannot consider it a "fraud" because the phenomenon is currently not regulated. In Appendix B we discuss this subject in depth.
7.1 Looking for 1-day Rug Pulls
We leverage our datasets to identify rug pulls systematically. Since we saw a considerable number of 1-day tokens and most of them are created serially, we narrow our investigation to the 332,265 in BSC and 25,180 in Ethereum 1-day tokens with a liquidity pool. Given the duration of these operations, we will refer to them as 1-day rug pulls. We analyze all the Events emitted by the liquidity pools, looking for all the pools that emitted only one Mint and one Burn event in which the address that performs the transaction burns at least 99% of the minted LP-tokens (we don’t use 100% since a small fraction of tokens might be stuck in the wallet due to rounding).
7.1.1 Estimating the gains of the operations
The simple operation, where the attacker does not swap in his liquidity pool, can be carried out by performing just four transactions: A transaction that creates the token, one that creates the liquidity pool, one to add the liquidity, and finally, the last transaction to remove the liquidity. These transactions can be performed individually, or they can be aggregated by leveraging a smart contract. Of course, we consider both cases when computing the fees. If the attacker performs swaps on the liquidity pool, we also consider the transaction fees paid for each swap.
@@ 7.1.1 Estimating the gains of the operations
The formula can be split into two components. The first part computes the gain in the case of the simple operation.The second formula takes into account the improved version of the operation, where the creator of the liquidity pool manipulates it by performing swaps operations. In this case, we remove from the gain $T_{in}$, that is the amount of tokens that the manipulator artificially adds to the liquidity. We also add to the gain $T_{out}$, the quantity of tokens that the manipulator removes from the liquidity pool before the final removal of the liquidity ($T_{out}$). Finally, we remove from the gain the fees used to perform the swap operations ($fee_{swap}$).
@@ 7.2 Results
After processing our data, we discover that an incredibly high number of liquidity pools are actually rug pulls. In BSC, 272,349 out of 332,265 (81.2%) of the considered liquidity pools have a rug pull pattern, while 21,742 out of 25,180 (86.3%) in Ethereum. This result shows that attackers use most of the 1-day tokens as disposable to carry out rug pulls.
These operations are arranged by 116,516 different addresses in BSC and 16,539 different addresses in Ethereum. As we can expect from the previous analyses, most of the token spammers that operate in BSC are linked to this kind of activity. Indeed, in BSC, 2,112 out of 4,231 (50%) token spammers performed at least one rug pull. Instead, in Ethereum, there are only 45 token spammers (0.3%) that have been involved in this activity. We find 115 addresses that perform more than 100 rug pulls in BSC, accounting for 19.1% of the operations, with the most active performing 16,102 operations. Instead, in Ethereum, we find only one address performing more than 100. Interestingly, combining the information in the BSC and Ethereum dataset, we find a token spammer that operated on both blockchains with the same address 5550x87605612492c74bA0037fFaef676c0f3f6958918. He performs five rug pulls on Ethereum and three on BSC.
Looking at the liquidity pools, we find that BNB (97.8% of the cases) is the token paired the most with the 1-day token. It is followed by USDT (0.67%) and BUSD (0.15%), two stablecoins pegged to the USD. Instead, Wrapped Ether is paired with all the 1-day tokens in almost all the liquidity pools with a rug pull in Ethereum. As the next step, we want to estimate the number of users that fall prey to such activities. To do so, we exclude the addresses that swap into liquidity pools they have created themselves from this analysis. We collect 251,250 different addresses in BSC and 57,552 in Ethereum that interact with at least one liquidity pool with a rug pull pattern. These addresses performed 2,903,022 swaps on the considered liquidity pools in BSC and 317,257 in Ethereum.
We divide the swaps into buy (1-day token) and sell operations. As we can expect, given the anatomy of the 1-day rug pull, we find that most of the operations are buy operations. More in detail, in BSC 2,286,056 (78.7%) are buy operations and 616,966 (21.3%) sell operations. In Ethereum, we find a very similar pattern, with 254,061 (80.1%) buy operations and 63,196 (19.9%) sell operations.
As final metric, we compute the average value of the swaps performed by the users. The average amount of swaps is almost identical for buy and sell operations on both the blockchains, with 0.01 BNB for BSC and 0.19 ETH for Ethereum’s liquidity pools. Interestingly, we notice a considerable difference in the average swap amount between the two blockchains. Indeed, the average swap is approximately \$3 on BSC and \$360 on Ethereum.
@@ 7.2.1 The gains
Before computing the gains of the attackers, we calculate the average price an attacker has to invest to arrange the operation. If the attacker does not perform any swap into the liquidity pool, the cost of the operation is on average 0.03 BNB in the case of BSC and 0.2 ETH for the Ethereum blockchain. Thus, the investment needed to perform such operations is low, even if it could vary substantially when the blockchains are overloaded. For instance, we found some rug pulls that reached the cost of 1.1 BNB or even 3.3 ETH. The base cost to arrange the operation is interesting because it represents a bound to the loss the attackers have to afford for each operation.
@@
We leverage our datasets to compute the gain of the operation using the formula 1 described in Sec. 7.1.1. We describe the 266,340 operations on BSC and the 21,594 on Ethereum in terms of successful and unsuccessful operations based on the operation’s net gain. In particular, we consider an operation successful if the net gain is strictly positive.
@@
Successful operations. Among the liquidity pools with a rug pull pattern, there are 104,404 (39.1%) operations in BSC and 13,368 (61.9%) in Ethereum closed with a profit for the attacker. A possible reason for the higher success rate of the rug pull on Ethereum could be that, as we saw, on average, users tend to invest more money. Indeed, on average, attracting only one investor is enough to cover the operation’s cost. To investigate what can affect the gains, we combine information on gains with those of the manipulations. When the creator of the liquidity pool does not perform any kind of manipulation, the net gain is, on average 0.11 BNB in BSC and 1.34 ETH in Ethereum. Operations carried out on liquidity pools that suffer wash-trading activity have an average gain of 0.25 BNB in BSC and 12 ETH in Ethereum, which is considerably higher than the previous case. Instead, we notice a negligible increase in gains in the case of pump operations with respect to the gains obtained by the liquidity pools without manipulation. Moreover, we find that both kinds of manipulation have no impact on the success rate. This show that operations that have wash trading are generally more profitable. However, the attacker has to perform several swaps, increasing its cost and loss in case of an unsuccessful operation.
@@
Unsuccessful operations. There are 161,936 (60.9%) liquidity pools in BSC and 8,226 (38.1%) in Ethereum, for which the attacker does not cover the transaction fees with the operations. For the 14% (21,122) of these liquidity pools of BSC and the 20% (1,506) of Ethereum, we notice that the operations were unsuccessful because nobody swapped into the liquidity pools. Considering the results we obtained, we can conjecture that the aim of the attackers is not to be successful every time but to arrange rug pulls serially and take profit in the long run. Indeed, the loss of an unsuccessful operation is minimal, and a streak of operations closed in loss can be covered with a single profitable operation.
Financial cost of 1-day rug pulls and comparison with other blockchain phenomena. In our study, we find that the number of 1-day rug pulls (21,594) and attackers (16,439) in Ethereum is significantly lower than in BSC (266,340 operations carried out by 117,110 rug pullers). Nonetheless, the total gain of Ethereum operations, around \$150 million, is remarkably higher than the gains of BSC operations, that amount to \$91 million. Moreover, the same trend holds when considering the volume of rug pull operations, which we define as the total value of BNB and ETH swapped. Here we find that Ethereum has a volume of \$772.5 million against the \$243.5 million of BSC. To gain insight into the magnitude of 1-day rug pull operations, we compare our metrics with popular blockchain shenanigans, like MEV and front-running. Tab. 4 in Appendix reports more relevant metrics collected from related works about operations carried out in Ethereum. As we can see, 1-day rug pull is the second type of operation by profit, generating slightly lower gains than Sandwich Attacks (\$174.34 million in accordance with Qin et al. [50]). Particularly interesting is the number of addresses that perform the operations. Indeed, in Ethereum, the number of attackers that performed 1-day rug pulls is almost five times the number of the Sandwich Attackers (the fraud with the higher number of attackers in our comparison). We believe the operations are performed by a large number of addresses due to their ease of execution. The reported numbers highlight that 1-day rug pulls is a significant phenomenon in the DeFi ecosystems that involve hundreds of thousands of malicious actors and move more than 1 billion USD.
7.2.2 A longitudinal view
Figure 4: The figure shows the number of rug pull operations (a), the initial liquidity added to each pool (b), and the gain for each operation over time. All the metrics are aggregated daily. The dashed vertical lines divide the three phases we identify.
Fig. 4 provides a longitudinal view of the daily number of rug pull operations (Fig. 4 a), the liquidity added (Fig. 4 b) and the gains (Fig. 4 c). Analyzing the trends of the chart, we identify three different phases, divided by the black dashed lines in the figure. In the first phase, we find the first spike of 1-day rug pulls in Ethereum. In the second phase, rug pulls start to increase in the BSC. However, the Ethereum gains are generally higher for the same invested liquidity. Finally, in the third phase, we see that BSC surpasses Ethereum in terms of liquidity added, number of operations, and gains. In the following, we describe in detail the three phases:
Phase 1: DeFi Summer. The first phase took place approximately from June 2020 to March 2021. At the beginning of this phase, we see an increase in the daily number of rug pull operations in Ethereum, with a peak of 179 daily operations in October 2020. Then, the number of operations steadily decrease until March 2021. We believe that the increase in the number of operations was bootstrapped by a phenomenon known in the crypto-community as DeFi Summer 2020 [42]. During this period, DeFi became extremely popular, and, as a result, the market capitalization and prices of several tokens soared [51]. This interest in DeFi attracted new users looking for investment opportunities, which may have triggered the increase in rug pull operations. Fig. 4 (b) shows that there is a significant amount of liquidity invested in these operations, on average \$37,941 (44 ETH), with an average gain of \$5,969 (5.65 ETH) (Fig. 4). Note that this phase involves only the Ethereum blockchain because the BSC was released in September 2020 and was not very popular yet.
Phase 2: Altcoin season. Fig. 4 (a) shows a second spike in the number of rug pulls from March 2021 to September 2021. In this case, the spike involves both Ethereum and the BSC, which reach a maximum peak of 195 and 2,309 daily operations. It is interesting to notice that the number of operations over time follows the same trend for Ethereum and BSC. For this reason, we believe an exogenous event caused this spike. Analyzing the events of that period, we believe this rise in the number of operations may be a so-called Altcoin Season. An Altcoin Season is a period in which Altcoins [^6] Altcoins [32] is a combination of the two words ”alternative” and ”coin”. The term is used to indicate all cryptocurrencies except Bitcoin. perform better than Bitcoin, significantly increasing their value. Previous study [38] shows that an Alt Season is marked by a drop of an indicator called Bitcoin dominance. This indicator measures the ratio between the market capitalization of Bitcoin to the total market capitalization of the entire cryptocurrency market. According to Coinmarketcap [18], in this period, the Bitcoin dominance decreased from 69% of January 2021 to 39% in May 2021. This market phase is frequently characterized by "Fear of missing out" (FOMO) [5], which makes investors more inclined to buy riskier tokens. For this reason, we believe investors have flocked to AMM markets to buy tokens, and rug pull operations skyrocketed. Fig. 4 (b) shows that the liquidity invested in these operations is higher for Ethereum, with an average of \$39,625 (50 ETH) against the \$5,624 (15 BNB) of BSC. Operations in Ethereum are also way more profitable, with an average gain of \$5,836 (6.3 ETH) against the \$48.4 (0.12 BNB) of BSC operations.
Phase 3: The overtaking of the BSC. The last phase goes from October 2021 to March 2022. In this phase, we find an interesting twist, as BSC surpasses Ethereum in terms of liquidity added and gains of rug pull operations. Indeed, in this phase, rug pulls in BSC have significantly more liquidity invested than in the past (56.9 BNB on average vs. 15.3 BNB of the previous phase) and higher gains (2.26 BNB on average vs. 0.12 BNB of the previous phase). For this reason, we can see in Fig. 4 that the total daily invested liquidity and gains in BSC are significantly higher than Ethereum and reached more than one million USD. In Sec. 11, we explore some possible reasons for this increase.
7.2.3 Tokens’ names